OpenStudy (anonymous):

Which of the point-slope equations below is correct for the line that passes through points (-1,-4) and (3,2)? Check all that apply.

OpenStudy (anonymous):

A.y-1=3/2(x+3) B.y+4=3/2(x-1) C.y+4=3/2(x+1) D.y-2=3/2(x-3) E.y+4=2/3(x+1)

OpenStudy (igreen):

https://www.desmos.com/calculator

OpenStudy (anonymous):

@iGreen can u help me

OpenStudy (igreen):

C and D are the same, and they both cross through points (-1,-4) and (3,2). I used https://www.desmos.com/calculator to graph them.

OpenStudy (igreen):

Welcome to Open Study! You can give answers by clicking 'Best Response'.

OpenStudy (anonymous):

so the answer is c and d

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

yes, that's correct. @RajonRondo9